# Knowledge-Informed Multi-Agent Trajectory Prediction at Signalized Intersections for Infrastructure-to-Everything

### v1

@@ Abstract
Multi-agent trajectory prediction at signalized intersections is crucial for developing efficient intelligent transportation systems and safe autonomous driving systems. Due to the complexity of intersection scenarios and the limitations of single-vehicle perception, the performance of vehicle-centric prediction methods has reached a plateau. Furthermore, most works underutilize critical intersection information, including traffic signals, and behavior patterns induced by road structures. Therefore, we propose a multi-agent trajectory prediction framework at signalized intersections dedicated to Infrastructure-to-Everything (I2XTraj). Our framework leverages dynamic graph attention to integrate knowledge from traffic signals and driving behaviors. A continuous signal-informed mechanism is proposed to adaptively process real-time traffic signals from infrastructure devices. Additionally, leveraging the prior knowledge of the intersection topology, we propose a driving strategy awareness mechanism to model the joint distribution of goal intentions and maneuvers. To the best of our knowledge, I2XTraj represents the first multi-agent trajectory prediction framework explicitly designed for infrastructure deployment, supplying subscribable prediction services to all vehicles at intersections. I2XTraj demonstrates state-of-the-art performance on both the Vehicle-to-Infrastructure dataset V2X-Seq and the aerial-view dataset SinD for signalized intersections. Quantitative evaluations show that our approach outperforms existing methods by more than 30% in both multi-agent and single-agent scenarios.
1 Introduction
Trajectory prediction is vital in enabling autonomous vehicles to navigate safely and efficientlyCao (2024). Signalized intersections, which account for a disproportionate number of traffic accidentsEkmekci et al. (2024), present particularly complex challenges for autonomous systems. Infrastructure-based trajectory prediction systems, enabled by Vehicle-to-Everything (V2X) communication, can provide centralized trajectory prediction services to all vehicles approaching intersections. This approach not only enhances autonomous driving capabilities to reduce the occurrence of accidents but also significantly improves V2X communication dependabilityZeng et al. (2024).
Many studies propose vehicle-centric trajectory prediction models for highwaysLiao et al. (2024a), urban roadsGu et al. (2021), and unsignalized intersectionsRowe et al. (2023). These models show promising results in various driving scenarios. However, few studies address trajectory prediction at signalized intersections, where traffic control introduces two critical challenges. On the one hand, vehicle behavior patterns at intersections differ significantly from other urban roads due to road topology and traffic controlSingh et al. (2022). This behavioral complexity requires a redundant prediction model to handle signalized intersection scenarios, placing a heavy computational burden on autonomous vehicles’ limited on-board resources. Infrastructure-deployed prediction models that broadcast predictive services offer a solution by centralizing these complex computations away from individual vehicles. On the other hand, due to inherent limitations in sensor coverage and field of view, individual on-board sensing systems cannot effectively capture the complete dynamic traffic environment at complex intersection scenarios. As Figure LABEL:fig:_introduction, while vehicles A and B are occluded from the autonomous vehicle’s (AV) field of view, they remain detectable by the roadside unit (RSU). With V2X technology, external information from other vehicles and infrastructures has been used to facilitate more accurate and reliable trajectory predictions than vehicle-based systemsRuan et al. (2023); Annunziata et al. (2024). Furthermore, the unidirectional transmission from infrastructure-to-vehicle(I2V) addresses privacy concerns associated with bidirectional vehicle-to-vehicle(V2V) and vehicle-to-infrastructure(V2I) communication. Infrastructure-to-Everything(I2X) trajectory prediction can facilitate communication device operations and enhance communication quality by enabling proactive network adjustments and optimized resource allocationSarkar and Taherkordi (2023); Wang et al. (2024a).
Therefore, we propose I2XTraj, an infrastructure-based trajectory prediction model for vehicles at signalized intersections. This approach also represents a novel paradigm for collaborative trajectory prediction. Our method performs trajectory prediction in a knowledge-driven manner with dynamic graph attention by leveraging real-time traffic information and prior behavior knowledge from infrastructure devices: (1) Real-time traffic signals. The continuous signal-informed mechanism is developed to encode signal sequences obtained directly from traffic light devices. Different types of intersections have varying rules and deployment of traffic lights. Our nonlinear continuous encoding function adaptively integrates color sequences and control directions. (2) Driving Behaviors at Intersections: The motion patterns and driving strategies of vehicles at intersections must conform to road topological structures while considering interaction events. Therefore, utilizing prior knowledge of intersection areas and vehicle driving states, we have designed a driving strategy awareness mechanism to estimate the distribution of such driving behaviors. The distribution represents a joint distribution of goal intentions and maneuvers for all vehicles. Based on the distribution, multimodal joint trajectory proposals are generated. (3) Multi-Agent Interaction: The spatial-temporal-mode attention is employed to model interaction relationships among agents, enabling fine-tuned adjustment of joint predicted trajectories. Additionally, an off-road loss is introduced to supervise the rationality of all trajectory modes. Finally, we validate our approach on the real-world dataset V2X-Seq and the signalized intersection drone dataset SinD. In single-infrastructure scenarios, our model provides reliable joint multi-agent trajectory predictions for all visible agents. In online collaborative scenarios, our method achieves more accurate single-agent prediction results for the target agent.
The principal contributions of this study are: (1) A novel trajectory prediction framework I2XTraj is proposed, which synthesizes a continuous signal-informed mechanism, driving strategy awareness mechanism, and spatial-temporal-mode attention in a knowledge-driven manner with dynamic graph attention. (2) To our knowledge, I2XTraj represents the first collaborative trajectory prediction paradigm specifically engineered for infrastructure in intersection scenarios. (3) Our method outperforms existing state-of-the-art methods by more than 30% in V2X-Seq and 15% in SinD datasets.
2 Related Work
This section reviews trajectory prediction at signalized intersections with traffic signals and unique driving behaviors.
@@ 2 Related Work
Trajectory Prediction for Signalized Intersections.
@@
Vehicle-based methods are constrained by their perceptual capabilities, lacking access to global trajectory information and real-time traffic control signals at intersections, leading to suboptimal performance. Consequently, cooperative trajectory prediction has been recognized as a promising solution to address these challengesWang et al. (2024b). In recent years, with the advancement of V2X technologies, several large-scale cooperative datasets that support prediction tasks have been releasedYu et al. (2023); Yan et al. (2023); Zhou et al. (2024). Researchers have established cooperative prediction paradigms at three stages: raw trajectory positionsRuan et al. (2023), cooperative featuresChen et al. (2024), and layered prediction resultsAnnunziata et al. (2024); Cao et al. (2024). However, these approaches necessitate frequent low-latency interactive communication between vehicles and infrastructure while simultaneously requiring precise trajectory prediction, imposing a substantial computational burden on vehicular platforms. Our method addresses the perceptual limitations by utilizing overhead roadside sensors, simplifying the complex task of joint prediction-communication optimization into a simple public information subscription task for vehicles.
Traffic Signals for Trajectory Prediction.
@@
Traffic signal states have been demonstrated to influence driving behavior at intersections significantlyPaul et al. (2022). In early studies, the states have been typically utilized in discrete forms. Discrete indexes have established dependencies for discontinuous vehicle motion behaviorsZhang et al. (2022). One-hot encoding has been used to predict vehicle intentionsCao et al. (2024) and simulate vehicle behaviorsWu et al. (2024). Dictionaries have embedded discrete states into high-dimensional spacesWei et al. (2024). However, these discrete methods neglect the temporal characteristics of signals. Furthermore, previous methods have been designed for standard orthogonal intersections, lacking generalizability across different intersection types. We propose a continuous signal-informed mechanism to encode continuous signals and adaptively process heterogeneous traffic control configurations.
Figure 2: The overall framework of our I2XTraj. Our architecture is an infrastructure-based method, which comprises three parts: (a) Knowledge-Informed Scene Encoding Module embeds agents’ historical states with traffic signal and map knowledge. (b) Driving Strategies-Aware Module generates strategy modes based on topological features and behavioral distributions to trajectory proposals. (c) Spatial-Temporal-Mode Attention Module spans the three dimensions to generate predicted scene trajectories.
Driving Behaviors for Trajectory Prediction.
@@
Effective mining of historical trajectory and traffic flow data can yield essential behavioral priors and patternsDing and Zhao (2023). Such knowledge is typically employed to design secondary tasks, encompassing goal-basedGu et al. (2021), maneuver-basedLiao et al. (2024b), and interaction-based Rowe et al. (2023) trajectory prediction. Infrastructure sensor devices enable the superior collection of behavioral distributions in intersection areas. Leveraging prior knowledge of driving behaviors and intersection topologies, the driving strategy awareness mechanism captures vehicles’ goal intentions and maneuvers, enabling multimodal prediction.
3 Problem Formulation
@@
The objective of multi-agent trajectory prediction is to forecast the future trajectories of all agents with historical states. The model can directly extract the state features of agents $H$ from the infrastructure within historical time steps $T_{h}$. Given that specific past states can lead to multiple possible futures, the goal of joint agents $N$ prediction is to predict trajectories for any agent $n$ under each possible scene $S^{k}$:
@@
4 Methodology
The proposed I2XTraj framework comprises three integral components, as illustrated in Figure 4: the Knowledge-Informed Scene Encoding Module, the Driving Strategy-Aware Module, and the Spatial-Temporal-Mode Attention Module. Each component is specifically engineered to address the unique challenges of infrastructure-based trajectory prediction at signalized intersections.
4.1 Knowledge-Informed Scene Encoding Module
Infrastructure-based trajectory prediction primarily requires encoding of intersection scenarios. Recent studies have demonstrated the remarkable effectiveness of Graph Neural Networks (GNNs) and relative spatiotemporal position encodingTang et al. (2024); Zhou et al. (2023). In the knowledge-informed scene encoding module, agent states, traffic light states, and HD maps are encoded as node attributes, while their relative spatio-temporal relationships are represented as edge attributes. This encoding method enables the framework to capture both the individual characteristics of scene elements and their intricate interconnections within the dynamic intersection environment.
Agent Encoder.
@@ 4.1 Knowledge-Informed Scene Encoding Module
Continuous Traffic Light Encoder.
I2XTraj deployed directly on infrastructure can conveniently utilize traffic light information, including its continuous digital signals, control directions, and traffic light locations. Inspired by the attention mechanismVaswani (2017), we encode the digital signals using a nonlinear continuous function similar to positional embedding:
@@
${A_{L}^{t,\ell}}=\text{MLP}(PE^{t,\ell})+\text{MLP}(d^{\ell}),$ (5)
where $d^{\ell}$ is the direction of traffic flow controlled by the traffic light. The composite feature representation encapsulates the traffic signal-governed intersection scenarios. This integrated representation captures both the instantaneous state of the intersection under traffic control and maintains the temporal context necessary for understanding the evolution of traffic patterns.
Map Encoder.
@@
${A_{M}^{m}}=\text{MLP}(l^{m},c^{m}),$ (6)
where $l^{m}$ represents the length of the centerlines, and $c^{m}$ denotes the attribute features of the lane segment. The positions and coordinates of the centerlines are embedded as edge features. LaneGCNLiang et al. (2020) is employed to extract the topological relationships between lanes, utilizing graph self-attention to capture the interaction features.
Graph Edge Encoder.
Similar to HPNetTang et al. (2024), we encode the relative spatio-temporal relationships between nodes as edge features in local polar coordinate systems. A two-layer MLP is employed to encode edge attributes:
$E_{e}=\text{MLP}(d^{e},\phi^{e},\psi^{e},\delta^{e}),$ (7)
where $d^{e}$, $\phi^{e}$, $\psi^{e}$ and $\delta^{e}$ represent spatial distance, the edge orientation in the reference frame, relative node orientation and temporal difference, respectively.
4.2 Driving Strategy-Aware Module

### v2

@@ Abstract
Multi-agent trajectory prediction at signalized intersections is crucial for developing efficient intelligent transportation systems and safe autonomous driving systems. Due to the complexity of intersection scenarios and the limitations of single-vehicle perception, the performance of vehicle-centric prediction methods has reached a plateau. In this paper, we introduce an Infrastructure-to-Everything (I2X) collaborative prediction scheme. In this scheme, roadside units (RSUs) independently forecast the future trajectories of all vehicles and transmit these predictions unidirectionally to subscribing vehicles. Building on this scheme, we propose I2XTraj, a dedicated infrastructure-based trajectory prediction model. I2XTraj leverages real-time traffic signal states, prior maneuver strategy knowledge, and multi-agent interactions to generate accurate, joint multi-modal trajectory prediction. First, a continuous signal-informed mechanism is proposed to adaptively process real-time traffic signals to guide trajectory proposal generation under varied intersection configurations. Second, a driving strategy awareness mechanism estimates the joint distribution of maneuver strategies by integrating spatial priors of intersection areas with dynamic vehicle states, enabling coverage of the full set of feasible maneuvers. Third, a spatial-temporal-mode attention network models multi-agent interactions to refine and adjust joint trajectory outputs. Finally, I2XTraj is evaluated on two real-world datasets of signalized intersections, the V2X-Seq and the SinD drone dataset. In both single-infrastructure and online collaborative scenarios, our model outperforms state-of-the-art methods by over 30% on V2X-Seq and 15% on SinD, demonstrating strong generalizability and robustness. The proposed I2X scheme and I2XTraj model offer a flexible solution for enhancing autonomous driving systems with infrastructure-enabled trajectory prediction.
Index Terms:
@@
Trajectory prediction, Signalized intersections, Multi-agent, Knowledge-driven, Infrastructure-to-Everything (I2X).
I Introduction
Trajectory prediction is vital in enabling autonomous vehicles to navigate safely and efficiently. Signalized intersections, which account for a disproportionate number of traffic accidents[1], present particularly complex challenges for autonomous systems. Infrastructure-based trajectory prediction systems, enabled by Vehicle-to-Everything (V2X) communication, can provide centralized trajectory prediction services to all vehicles approaching intersections. This approach not only enhances autonomous driving capabilities to reduce the occurrence of accidents but also significantly improves V2X communication dependability[2].
Many studies propose vehicle-centric trajectory prediction models for highways[3], urban roads[4], and unsignalized intersections[5]. These models show promising results in various driving scenarios. However, few studies address trajectory prediction at signalized intersections. A prototypical signalized intersection scenario is illustrated in Fig. LABEL:fig:_introduction, where the roadside unit (RSU) maintains full visibility of all vehicles. The regulation of traffic signal for vehicular flow imposes deterministic constraints on future trajectory prediction, whereas the multi‐branch characteristics of intersection lanes introduce uncertainty to maneuvering strategies available to vehicles.
Traffic signal information and prior knowledge of maneuvering strategies are not readily obtainable by on-board autonomous driving systems in isolation. With V2X technology, external information from other vehicles and infrastructures has been used to facilitate more accurate and reliable trajectory predictions than vehicle-based systems[6, 7]. These data sourced from infrastructure may be fused or exchanged at successive stages of an autonomous driving pipeline[8]. At the earliest stage, raw image or point‐cloud data are transmitted. Later stages exchange higher-level information, culminating in predicted vehicle trajectories. As processing depth increases, the volume of data requiring transmission between the autonomous vehicle (AV) and roadside units (RSUs) decreases. Consequently, the demands on temporal synchronization and communication latency are progressively reduced[9, 10].
In prediction‐mediated cooperative schemes, information exchange between vehicle and infrastructure typically proceeds through three stages, including historical trajectories[6], trajectory features[11, 12, 13], and future trajectories[7]. These cooperative schemes typically demand a high degree of coupling between the on-board model and roadside information, constraining the design of autonomous driving systems. To overcome this limitation, we propose an I2X collaborative prediction scheme, which forecasts all intersection vehicle trajectories independently within the infrastructure, then unidirectionally transmits the future trajectories to the vehicles. The roadside-predicted trajectories can serve as input to any type of autonomous driving system, thereby endowing the on-board subsystem with considerable design flexibility.
Within the I2X collaborative prediction scheme, trajectory prediction models deployed on infrastructure play a critical role. On the one hand, the inherent characteristics of such infrastructure information also introduce unique challenges. In particular, the infrastructure is unable to distinguish between autonomous and conventional vehicles traversing an intersection, and may observe multiple autonomous vehicles simultaneously. Consequently, the prediction model must be capable of forecasting the trajectories of all vehicles at the intersection, thereby enabling autonomous vehicles to utilize the provided prediction outputs selectively. On the other hand, the supplementary information provided by the infrastructure offers actionable insights for multi-agent trajectory prediction, particularly in scenarios involving a high density of traffic participants. This information comprises real-time traffic signal states and prior knowledge of intersection vehicle maneuvering strategies. By modeling the dependence of vehicle trajectories on traffic control signals, the prediction model can reduce uncertainty in forecasts. Moreover, incorporating prior maneuver strategies enables the extraction of latent driving intentions, thereby covering the full set of potential feasible trajectories.
Therefore, we propose I2XTraj, an infrastructure-based trajectory prediction model for all vehicles at signalized intersections. Our method performs trajectory prediction in a knowledge-driven manner with dynamic graph attention by leveraging real-time traffic information and prior maneuver strategy knowledge from infrastructure devices. Firstly, the continuous signal-informed mechanism is designed to represent signal sequences obtained directly from traffic light devices. Different types of intersections have varying rules and deployment of traffic lights. Our nonlinear continuous encoding function adaptively integrates color sequences and control directions to guide the generation of trajectory proposals. Secondly, the driving strategy awareness mechanism estimates the joint distribution of maneuver strategies, utilizing prior knowledge of intersection areas and vehicle driving states. Based on the distribution, multi-modal joint trajectory proposals are generated. Thirdly, the spatial-temporal-mode attention is employed to model interaction relationships among agents, enabling fine-tuned adjustment of joint predicted trajectories. Finally, we validate our approach on the real-world dataset V2X-Seq and the signalized intersection drone dataset SinD. In single-infrastructure scenarios, our model provides reliable joint multi-agent trajectory predictions for all visible agents. In online collaborative scenarios, our method achieves more accurate single-agent prediction results for the target agent.
The principal contributions of this study are:
(1) The I2X collaborative prediction scheme is proposed, wherein the infrastructure predicts the trajectories of all vehicles at signalized intersections independently and transmits them to subscribing vehicles unidirectionally.
(2) A dedicated infrastructure knowledge-informed trajectory prediction model in the I2X scheme, I2XTraj, is designed to accurately forecast the joint future trajectories of all vehicles at intersections, leveraging traffic signal states, maneuver strategy knowledge, and interaction information effectively.
(3) Our method outperforms existing state-of-the-art methods by more than 30% in the real-world V2I dataset V2XSeq and 15% in the drone dataset at signalized intersection SinD. The generalizability across different intersections and the robustness are also validated.
The remaining sections of this paper are organized as follows. Section II reviews the related work. Section III outlines the cooperative scenario and problem formulation. Section IV details the methodology of our trajectory prediction method. Section V thoroughly analyzes the experiment’s results and evaluates the model’s performance. Finally, we summarizes the entire work in Section VI.
II Related Work
@@
Most existing trajectory prediction studies focus on general urban environments[14, 15] or unsignalized intersections[16, 17], where the contextual information is predominantly obtained through vehicle-based sensors. Vehicle-based methods are constrained by their perceptual capabilities, lacking access to global trajectory information and real-time traffic control signals at intersections, leading to suboptimal performance. Consequently, cooperative trajectory prediction has been recognized as a promising solution to address these challenges[18]. In recent years, with the advancement of V2X technologies, several large-scale cooperative datasets that support prediction tasks have been released[19, 20, 21].
This section reviews cooperative trajectory prediction at signalized intersections with traffic signals and unique driving maneuvers.
II-A Cooperative Trajectory Prediction
The existing V2X research predominantly focuses on cooperative perception[22, 23, 24], while collaboration approaches leveraging prediction as a nexus have progressively garnered attention among researchers, due to their effective utilization of multi-source information and interaction characteristics. Researchers have established cooperative prediction paradigms at three stages: historical trajectory positions[6], cooperative features[11, 12, 13], and layered prediction results[7, 25]. However, these approaches rely on a highly coupled vehicle–infrastructure cooperative system and ultra-low communication latency, which imposes significant constraints on the development of autonomous driving models. Our method addresses the perceptual limitations by utilizing overhead roadside sensors, simplifying the complex task of joint prediction-communication optimization into a simple public information subscription task for vehicles.
II-B Traffic Signal Representations for Trajectory Prediction
@@
Traffic signal states have been demonstrated to influence driving maneuvers at intersections significantly[26]. In early studies, the states have been typically utilized in discrete forms. Discrete indexes have established dependencies for discontinuous vehicle motion behaviors[27]. One-hot encoding has been used to predict vehicle intentions[25] and simulate vehicle behaviors[28]. Dictionaries have embedded discrete states into high-dimensional spaces[29]. Bayesian networks have also been employed to infer discrete intentions[30]. However, these discrete methods neglect the temporal characteristics of signals. Furthermore, previous methods have been designed for standard orthogonal intersections, lacking generalizability across different intersection types. We propose a continuous signal-informed mechanism adaptively transforms raw traffic light color sequences and control directions into a nonlinear representation that guides trajectory proposal generation under varied intersection configurations.
Figure 2: A comparison between the Vehicle-Infrastructure Cooperation (VIC) pipeline and our proposed Infrastructure-to-Everything (I2X) cooperative prediction pipeline. (a) The VIC pipeline requires the autonomous vehicle (Ego AV) and infrastructure to be equipped with a unified system, enabling bidirectional communication at one or multiple stages. (b) The I2X cooperative pipeline does not require vehicles to send information to the infrastructure; instead, the infrastructure unidirectionally provides future trajectories to any type of autonomous driving system.
II-C Trajectory Prediction with Knowledge Priors
@@
Effective mining of historical trajectory and traffic flow data can yield essential behavioral priors and patterns[31]. Such knowledge is typically employed to design secondary tasks, encompassing goal-based[4], maneuver-based[32], and interaction-based[5] trajectory prediction. The incorporation of hard kinematic constraints and soft constraints derived from traffic regulations enhances both the reliability and interpretability of the prediction[33]. Infrastructure sensor devices enable the superior collection of maneuver distributions in intersection areas. Leveraging prior knowledge of driving maneuver strategies and intersection topologies, the driving strategy awareness mechanism captures intentions of vehicles, enabling multi-modal prediction to cover all potential maneuvers.
III Cooperative Scenario and Problem Formulation
III-A Infrastructure-to-Everything Cooperative Scenario
In existing collaborative trajectory prediction approaches, Vehicle-Infrastructure Cooperation (VIC) methods are commonly employed, as illustrated in Fig. 2(a). VIC methodologies can be categorized into two types[19]: online VIC and offline VIC. Online VIC refers to the process where vehicles and infrastructure synchronously conduct real-time data transmission and fusion during collaboration. Typical online VIC approaches can be categorized based on the levels at which data fusion is performed. PP-VIC [19] supplies the ego vehicle with infrastructure-side data in a frame-by-frame manner over the historical time horizon. The feature fusion approach [6] leverages motion and interaction features to enable cooperative trajectory feature fusion. The layered approach [7] exploits synergistic collaboration between infrastructure and vehicle layers to enhance prediction performance. The offline VIC approach first pre-trains models on infrastructure data before deploying them to vehicles for further training. However, both offline and online methods require highly coupled or centralized training between vehicles and infrastructure, which is challenging to implement in practical deployments. This is because real-world scenarios cannot mandate all autonomous vehicles to be equipped with identical cooperative systems or autonomous driving models, nor can they require vehicles to share their real-time data with infrastructure actively. Therefore, we propose the I2X prediction collaboration scheme, illustrated in Fig. 2(b), which involves deploying specialized prediction models for infrastructure and transmitting the future trajectories to all vehicles.
Since it is impossible to distinguish whether vehicles are equipped with autonomous driving systems from an infrastructure perspective, it cannot be applied that conventional methods of selecting target agents surrounding the Ego AV on RSUs. This issue has never been addressed in any existing VIC research. Consequently, specialized models deployed on infrastructure need to predict joint future trajectories for all vehicles in the scene and transmit them collectively to subscribing vehicles, allowing these vehicles to select agents of interest independently. Depending on the level of traffic congestion, I2XTraj is required to predict from a few to well over one hundred agents.
This collaborative approach provides considerable flexibility to autonomous vehicles, enabling any type of autonomous driving system to utilize these additional future trajectories from a global perspective. For planning and control systems, these future trajectories can either serve directly as part of the planning input or guide perception or prediction processes. Any other heterogeneous autonomous driving systems can also subscribe to these future trajectories as supplementary information.
III-B Problem Definition
Figure 3: The overall framework of our I2XTraj. Our architecture is an infrastructure-based method, which comprises three parts: (a) Knowledge-Informed Scene Encoding Module embeds agents’ historical states with traffic signal and map knowledge. (b) Driving Strategy-Aware Module generates strategy modes based on topological features and maneuver strategies distributions to trajectory proposals. (c) Spatial-Temporal-Mode Attention Module spans the three dimensions to generate predicted scene trajectories.
@@ III-B Problem Definition
The objective of multi-agent trajectory prediction is to forecast the future trajectories of all agents with historical states. The model can directly extract the state features of agents $\mathrm{H}$ from the infrastructure within historical time steps $T_{h}$. Given that specific past states can lead to multiple possible futures, the maneuver of joint agents $N$ prediction is to predict trajectories for any agent $n$ under each possible scene ${S}^{k}$:
@@
IV Methodology
The proposed I2XTraj framework comprises three integral components, as illustrated in Fig. 3: the Knowledge-Informed Scene Encoding Module, the Driving Strategy-Aware Module, and the Spatial-Temporal-Mode Attention Module. Each component is specifically engineered to address the unique challenges of infrastructure-based trajectory prediction at signalized intersections.
IV-A Knowledge-Informed Scene Encoding Module
Infrastructure-based trajectory prediction primarily requires encoding of intersection scenarios. Recent studies have demonstrated the remarkable effectiveness of Graph Neural Networks (GNNs) and relative spatiotemporal position encoding[34, 35]. In the knowledge-informed scene encoding module, agent states, traffic light states, and HD maps are encoded as node attributes, while their relative spatio-temporal relationships are represented as edge attributes. This encoding method enables the framework to capture both the individual characteristics of scene elements and their intricate interconnections within the dynamic intersection environment.
Agent Encoder
@@ IV-A Knowledge-Informed Scene Encoding Module
Continuous Traffic Light Encoder
I2XTraj deployed directly on infrastructure can conveniently utilize traffic light information, including its continuous digital signals, control directions, and traffic light locations. As shown in Fig. 4, we encode the digital signals using a nonlinear continuous function similar to positional embedding:
@@
Figure 4: Schematic illustration of the continuous signal-informed mechanism.
@@
${{A}_{L}^{t,\ell}}=\text{MLP}({PE}^{t,\ell})+\text{MLP}({d}_{L}),$ (5)
where $d_{L}$ is the direction of traffic flow controlled by the traffic light. The composite feature representation encapsulates the traffic signal-governed intersection scenarios. This integrated representation captures both the instantaneous state of the intersection under traffic control and maintains the temporal context necessary for understanding the evolution of traffic patterns.
Map Encoder
@@
${{A}_{M}}=\text{MLP}({l}_{M},{c}_{M}),$ (6)
where $l_{M}$ represents the length of the centerlines, and ${c}_{M}$ denotes the attribute features of the lane segment. The attribute features include lane center position, lane heading, is intersection, turning sign, is controlled by traffic lights, centerline positions, and centerline heading. All of these features come from the HD map automatically. The positions and coordinates of the centerlines are embedded as edge features. LaneGCN[36] is employed to extract the topological relationships between lanes, utilizing graph self-attention to capture the interaction features.
Graph Edge Encoder
Similar to HPNet[34], we encode the relative spatio-temporal relationships between nodes as edge features in local polar coordinate systems. A two-layer MLP is employed to encode edge features:
${E}_{e}=\text{MLP}({d}_{e},{\phi}_{e},{\psi}_{e},{\delta}_{e}),$ (7)
where ${d}_{e}$, ${\phi}_{e}$, ${\psi}_{e}$ and ${\delta}_{e}$ represent spatial distance, the edge orientation in the reference frame, relative node orientation and temporal difference, respectively.
Figure 5: Schematic illustration of the driving awareness mechanism. The red circular area indicates the range of the stop maneuvers. The blue area represents the range of U-turn maneuvers. The yellow area corresponds to the range of left-turning maneuvers, while the green area denotes the range of maneuvers moving straight. The orange area indicates the range of right-turning maneuvers.
IV-B Driving Strategy-Aware Module